# The Geometry of Multi-Task Grokking: Transverse Instability, Superposition, and Weight Decay Phase Structure

### v1

@@ Abstract
We find five consistent phenomena. (1) Staggered grokking order: multiplication generalizes first, followed by squaring and then addition, with consistent delays across seeds. (2) Universal integrability: optimization trajectories remain confined to an empirically invariant low-dimensional execution manifold, with commutator defects orthogonal to this manifold and defect onset reliably preceding generalization ($42/42$ conditions). (3) Weight decay phase structure: grokking timescale, curvature depth, reconstruction threshold, and defect lead covary systematically with weight decay, revealing distinct dynamical regimes and a sharp no-decay failure mode. (4) Holographic incompressibility: although final solutions occupy only 4–8 principal trajectory directions, they are distributed across full-rank weights and are destroyed by minimal perturbations; neither per-layer SVD, magnitude pruning, nor uniform scaling preserves performance. (5) Transverse fragility and redundancy: removing less than 10% of orthogonal gradient components eliminates grokking, yet dual-task models exhibit partial recovery under extreme deletion, suggesting redundant center manifolds enabled by overparameterization.
@@ 1 Introduction
Section 2 describes the experimental setup. Sections 3, 4, 5, 6, 7 and 8 present empirical results across six domains: training dynamics, manifold structure, integrability, Hessian analysis, the weight decay phase diagram, and incompressibility. Section 9 synthesizes these findings into a unified account of superposition dynamics in parameter space. Section 10 develops a theory of overparameterization as manifold redundancy, grounded in the transverse ablation results. Section 11 connects our findings to broader themes, discusses limitations, and proposes interpretive frameworks.
@@ 9.2 Sharp Reconstruction Phase Transition
Reconstruction accuracy as a function of $k$ exhibits a sharp phase transition (cf. Montanari and Wang, 2026) (see Figures 29 and 30 and Tables 12 and 13 in the Appendix for detailed per-$k$ accuracy breakdowns):
@@ 9.4 Role of Regularization: Compression Pressure
Dual-task models systematically require more PCA components than tri-task models under comparable weight decay (Tables 8, 9 and 19). For $\lambda\in[0.1,1.0]$, tri-task models typically require $k^{*}\approx 4$66 components, while dual-task models require $k^{*}\approx 5$88. Lower weight decay further increases $k^{*}$ in both settings. This indicates that solution dimensionality increases when task constraints are relaxed.
@@
Figure 19: Constraint-induced compression. (a) Dual-task models require $k^{*}=5$$10$ PCA directions depending on WD, with the orthogonal complement carrying 0.2–0.8% of trajectory variance. (b) Tri-task models require fewer components ($k^{*}=3$88), consistent with stronger superposition pressure from additional task constraints. See Tables 13 and 12 for per-$k$ accuracy breakdowns.
@@ 10.1 Transverse Gradient Ablation
where $\alpha\in[0,1]$ controls deletion strength. We then measured grokking time under varying $\alpha$ (see Tables 10 and 11 for full results, and Tables 14 and 15 in the Appendix for additional detail).
@@ Appendix A Additional Figures
Figure 29: Detailed comparison of reconstruction threshold $k^{*}$ between dual-task and tri-task settings (companion to Figure 19 in the main text). (a) Dual-task models require more PCA directions; the chance-to-perfect transition is visible as a sigmoid-like jump. (b) Tri-task models exhibit the same sharp transition but shifted leftward to lower $k$, indicating stronger constraint-induced compression. The systematic $k^{*}_{\text{dual}}>k^{*}_{\text{tri}}$ ordering holds across all five WD values and three random seeds (see Tables 13 and 12 for numerical details).

### v3

@@ Abstract
[…]
@@ 1 Introduction
Section 2 describes the experimental setup. Sections 3, 4, 5, 6, 7 and 8 present empirical results across six domains: training dynamics, manifold structure, integrability, Hessian analysis, the weight decay phase diagram, and incompressibility. Section 9 analyzes the spectral geometry of the attention operator transition, revealing universal phase trajectories and layer asymmetry in instability magnitude. Section 10 synthesizes these findings into a unified account of superposition dynamics in parameter space. Section 11 develops a theory of overparameterization as manifold redundancy, grounded in the transverse ablation results. Section 12 connects our findings to broader themes, discusses limitations, and proposes interpretive frameworks.
@@ 8.3 Orthogonal Fragility
9 Spectral Geometry of the Attention Operator Transition
The preceding sections characterize grokking through trajectory PCA, commutator defects, and Hessian curvature. We now ask: what happens inside the attention operators during the commutator transition? To answer this, we analyze the singular value spectrum of the attention weight matrices $W_{Q}$ and $W_{K}$ at each training checkpoint for both dual-task and tri-task models, complementing the single-task spectral analysis of a companion study (Xu, 2026b).
Geometric picture.
These observations suggest that grokking corresponds to a geometric phase transition in the attention operators. Early in training, the dominant singular directions of $W_{Q}$ and $W_{K}$ are nearly degenerate and the operators exhibit strong non-commutativity. As training proceeds, a symmetry-breaking event occurs in which one singular direction becomes dominant, the commutator defect collapses, and the attention operators align into a shared eigenbasis. Generalization emerges only after this alignment is achieved. The spectral gap, the matrix commutator, and grokking are thus linked by a single dynamical narrative: degeneracy breeds instability, symmetry breaking resolves it, and generalization follows.
9.1 Spectral Symmetry Breaking Precedes Grokking
Define the leading spectral gap of the query matrix at layer $\ell$:
$g_{12}^{(\ell)}=\sigma_{1}(W_{Q}^{(\ell)})-\sigma_{2}(W_{Q}^{(\ell)}).$ (2)
In every grokking run we examine—dual-task (3 seeds) and tri-task (3 seeds)—the spectral gap opens during or immediately before commutator collapse (Figure 19). The attention operator transitions from a near-degenerate regime ($\sigma_{1}\approx\sigma_{2}$) to a rank-1-dominated regime ($\sigma_{1}\gg\sigma_{2}$), and this transition is a necessary precursor to generalization. Memorizing (no–weight-decay) runs show no such symmetry-breaking event: their spectral gaps remain small and the commutator wanders diffusively (Figure 20).
9.2 Grokking Follows a Universal Phase Trajectory
In the $(\sigma_{1}-\sigma_{2},\;\left\|[W_{Q},W_{K}]\right\|_{F})$ phase plane, every grokking run traces a characteristic loop through three regimes:
  1. 1.
    Competition: small spectral gap, rising commutator;
  2. 2.
    Instability: spectral gap opening, commutator at peak;
  3. 3.
    Alignment: large spectral gap, collapsing commutator; generalization appears.
This loop structure is observed in all single-task operations (companion study), in dual-task arithmetic, and in tri-task arithmetic (Figures 19 and 20). Memorizing runs exhibit no loop structure: their trajectories are diffuse random walks with substantially higher commutator values and no directed escape. The universality of the loop suggests that grokking reflects a generic geometric transition in the attention operator rather than a task-specific phenomenon.
9.3 Multi-Task Grokking Occurs Along a Shared Trajectory
In the tri-task setting, all three tasks grok at different training steps, yet the shared-trunk attention operators trace a single trajectory in phase space. Each task generalizes at a different position along this shared path (Figure 19):
$\underbrace{\text{comm.\ peak}}_{\text{step 17{,}600}}\;\to\;\underbrace{\text{grok:mul}}_{\text{step 17{,}900}}\;\to\;\underbrace{\text{grok:}a^{2}\!+\!b^{2}}_{\text{step 24{,}800}}\;\to\;\underbrace{\text{grok:add}}_{\text{step 30{,}500}}$ (3)
In the dual-task setting, the two tasks grok within 100–400 steps of each other (Table 1), consistent with the shorter trajectory and the near-simultaneity of operator alignment for two tasks. This result implies that multi-task grokking does not require separate per-task circuits or independent phase transitions. Instead, training reorganizes the shared geometric structure of the attention operators, and individual tasks become solvable as the trajectory passes through their respective alignment thresholds. The commutator transition is a property of the trunk, not of any individual task head.
Figure 19: Phase portrait for tri-task arithmetic (seed 42, layer 0) in the $(\sigma_{1}-\sigma_{2},\;\left\|[W_{Q},W_{K}]\right\|_{F})$ plane. Left: colored by training step. Right: colored by mean test accuracy across the three tasks. The three tasks grok at different positions along a single shared trajectory: multiplication (step 17,900) near the commutator peak, $a^{2}\!+\!b^{2}$ (step 24,800) during descent, and addition (step 30,500) in the alignment region. The competition–instability–alignment structure is preserved from the single-task case.
Figure 20: Grokking vs. memorizing phase portraits for dual-task arithmetic (modular addition + multiplication, seed 42, layer 0). Left: grokking run ($\mathrm{wd}=1$) shows the characteristic competition–instability–alignment loop, with both tasks grokking near the transition to alignment. Right: memorizing run ($\mathrm{wd}=0$) shows a diffuse random walk with $\sim\!3\times$ higher commutator values and no directed escape.
9.4 Layer Asymmetry in Instability Magnitude
We decompose the matrix commutator by layer and examine the peak magnitude of $\left\|[W_{Q}^{(\ell)},W_{K}^{(\ell)}]\right\|_{F}$ during training. Across all conditions tested (Table 12), the bottom layer (Layer 0) develops substantially stronger non-commutativity than the top layer (Layer 1):
$\max_{t}\left\|[W_{Q}^{(0)},W_{K}^{(0)}]\right\|_{F}\;>\;\max_{t}\left\|[W_{Q}^{(1)},W_{K}^{(1)}]\right\|_{F},\qquad\text{ratio}\approx 1.3\text{--}2.4\times.$ (4)
This amplitude asymmetry is robust across all seeds and datasets; the peak timing order between layers is not consistent across seeds and should not be regarded as an invariant.
[…]
Dataset Seed L0 peak L1 peak Ratio
[…]
The per-layer phase portrait (Figure 21) provides a geometric interpretation: Layer 0 traces a large loop in the $(\sigma_{1}-\sigma_{2},\;\left\|[W_{Q},W_{K}]\right\|_{F})$ plane, extending to high commutator values and wide spectral gaps before collapsing near the grokking transition. Layer 1 traces a smaller loop that peaks earlier and collapses sooner. The bottom layer bears the dominant geometric instability during grokking, while the top layer stabilizes first. In the multi-task setting, this asymmetry is amplified: the ratio increases from $\sim\!1.1\times$ (single-task) to $\sim\!1.7$$2.4\times$ (dual- and tri-task), suggesting that the additional representational demands of multiple tasks are absorbed primarily by the lower layer.
Figure 21: Layer-wise phase portrait overlay for tri-task arithmetic (seed 42). Layer 0 (blue) traces a large loop reaching peak commutator $\approx 24.7$ at step 17,600, while Layer 1 (orange) remains below $\approx 11.3$. The bottom layer bears the dominant geometric instability. The tri-task amplifies the asymmetry relative to the single-task case ($2.2\times$ vs. $1.1\times$).
9.5 Gram Matrix Eigenvalue Gap Dynamics
The weight-matrix spectral analysis above examines individual attention operators $W_{Q}$ and $W_{K}$. A complementary view comes from the rolling-window Gram matrix $\boldsymbol{G}(t)=\boldsymbol{X}(t)\boldsymbol{X}(t)^{\top}$, formed from $W$ consecutive parameter updates $\boldsymbol{\delta}_{s}=\boldsymbol{\theta}_{s+1}-\boldsymbol{\theta}_{s}$ (all attention weights flattened into $\mathbb{R}^{p}$), which captures the spectral structure of the optimisation trajectory itself (Xu, 2026d). Three quantities are extracted via the SVD of the $W\times p$ update matrix $\boldsymbol{X}$ with singular values $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{W}$:
  • (i)
    The signal rank $k^{*}=\arg\max_{j}\,\omega_{j}\cdot(\sigma_{j}/\sigma_{j+1})$, where $\omega_{j}=\sigma_{j}/\sum_{i}\sigma_{i}$ weights the ratio by spectral mass at position $j$, suppressing spurious tail ratios outside the extreme aspect ratio regime.
  • (ii)
    The gap ratio $R=\sigma_{k^{*}}/\sigma_{k^{*}+1}$, measuring how sharply the signal separates from sub-dominant modes.
  • (iii)
    The sub-leading eigenvalue gap $g_{23}=\sigma_{2}^{2}-\sigma_{3}^{2}$, tracking the separation between the second and third eigenvalues of $\boldsymbol{G}$.
We compute all three quantities with $W=10$ across 18 grokking runs (12 single-task from 4 operations $\times$ 3 seeds, plus 3 dual-task and 3 tri-task, all with $\lambda>0$) and 18 matched controls ($\lambda=0$). Table 13 reports the full results.
Table 13: Rolling-window Gram matrix spectral analysis ($W=10$). $g_{23}^{\mathrm{early}}$: peak $g_{23}$ before grokking; $g_{23}^{\mathrm{grok}}$: $g_{23}$ at the grokking step; Decline $g_{23}^{\mathrm{early}}/g_{23}^{\mathrm{grok}}$; $R$: gap ratio at signal rank $k^{*}$. Single-task operations that do not grok within $2\times 10^{5}$ steps (x2_xy_y2, x3_xy) are omitted.
Setting $\boldsymbol{\lambda}$ Grok step $g_{23}^{\mathrm{early}}$ $g_{23}^{\mathrm{grok}}$ Decline $R_{\mathrm{early}}$ $k^{*}_{\mathrm{term}}$ Decl.
Single-task (mean $\pm$ std over 3 seeds)
add 1.0 2500–3100 $14.6\pm 0.2$ $6\pm 0.4 50×50\times$ $1.40\pm 0.05$ 1
0.0 $33.0\pm 9.3$ $2.72\pm 0.22$
mul 1.0 2600–3000 $14.7\pm 0.1$ $0.39\pm 0.2$ $49\times$ $1.41\pm 0.05$ 1
0.0 $23.6\pm 10.7$ $2.85\pm 0.11$
sub 1.0 3300–3700 $14.4\pm 0.1$ $0.48\pm 0.2$ $35\times$ $1.31\pm 0.02$ 1
0.0 $33.2\pm 2.8$ $2.62\pm 0.23$
$a^{2}{+}b^{2}$ 1.0 2000–2500 $18.9\pm 0.5$ $0.94\pm 0.5$ $28\times$ $1.47\pm 0.08$ 1
0.0 $28.3\pm 13.1$ $3.12\pm 0.64$
Dual-task: add + mul
1.0 9700–16100 $30.8\pm 0.2$ $2.94\pm 0.8$ $11\times$ $1.35\pm 0.10$ 1
0.0 $18.6\pm 4.0$ $1.96\pm 0.35$
Tri-task: add + mul + sub
1.0 13100–26400 $31.3\pm 0.3$ $0.53\pm 0.2$ $63\times$ $1.41\pm 0.11$ 1
0.0 $32.3\pm 0.7$ $1.77\pm 0.31$
Aggregate
All $\lambda>0$ $39\times$ mean $1.39\pm 0.08$ 1 (15/18) 18/18
All $\lambda=0$ $2.51\pm 0.56$ 1/18
[…]
The $\lambda=0$ controls display the predicted spectral signature of the grokking-absent regime (Xu, 2026d, Remark 12.22): $g_{23}$ remains high throughout all training steps with no decline, confirming that without the curvature floor $\omega$ the spectral symmetry-breaking transition never initiates.
The Gram matrix analysis thus provides a complementary perspective to the weight-matrix SVD of the preceding subsections. While the weight-matrix spectral gap $g_{12}=\sigma_{1}(W_{Q})-\sigma_{2}(W_{Q})$ tracks the internal structure of individual attention operators, the Gram matrix $g_{23}$ tracks the coherence of the optimisation trajectory itself. Both converge on the same mechanistic picture: grokking is a spectral phase transition in which concentrated update dynamics give way to distributed circuit formation, with weight decay providing the driving force.
@@ 10.2 Sharp Reconstruction Phase Transition
Reconstruction accuracy as a function of $k$ exhibits a sharp phase transition (cf. Montanari and Wang, 2026) (see Figures 32 and 33 and Tables 14 and 15 in the Appendix for detailed per-$k$ accuracy breakdowns):
@@ 10.4 Role of Regularization: Compression Pressure
Dual-task models systematically require more PCA components than tri-task models under comparable weight decay (Tables 8, 9 and 22). For $\lambda\in[0.1,1.0]$, tri-task models typically require $k^{*}\approx 4$66 components, while dual-task models require $k^{*}\approx 5$88. Lower weight decay further increases $k^{*}$ in both settings. This indicates that solution dimensionality increases when task constraints are relaxed.
@@
Figure 22: Constraint-induced compression. (a) Dual-task models require $k^{*}=5$$10$ PCA directions depending on WD, with the orthogonal complement carrying 0.2–0.8% of trajectory variance. (b) Tri-task models require fewer components ($k^{*}=3$88), consistent with stronger superposition pressure from additional task constraints. See Tables 15 and 14 for per-$k$ accuracy breakdowns.
@@ 11.1 Transverse Gradient Ablation
where $\alpha\in[0,1]$ controls deletion strength. We then measured grokking time under varying $\alpha$ (see Tables 10 and 11 for full results, and Tables 16 and 17 in the Appendix for additional detail).
@@ 12.3 Weight Decay as a Phase Parameter
The intra-signal gap framework (Xu, 2026d) provides a spectral account of this mechanism: weight decay promotes the spectral gap opening at $k^{*}=1$ that stabilises the dominant attention direction, while no-weight-decay runs remain in the degenerate regime with no gap dynamics and no generalisation (see Section 9.5 for the full quantitative analysis).
@@ Appendix A Additional Figures
Figure 32: Detailed comparison of reconstruction threshold $k^{*}$ between dual-task and tri-task settings (companion to Figure 22 in the main text). (a) Dual-task models require more PCA directions; the chance-to-perfect transition is visible as a sigmoid-like jump. (b) Tri-task models exhibit the same sharp transition but shifted leftward to lower $k$, indicating stronger constraint-induced compression. The systematic $k^{*}_{\text{dual}}>k^{*}_{\text{tri}}$ ordering holds across all five WD values and three random seeds (see Tables 15 and 14 for numerical details).